OpenStudy (anonymous):

√((27w^10)/((w^8x^5z^4)))√((x^9z^4)/(3))

OpenStudy (anonymous):

\[\sqrt{\frac{ 27w^10 }{ w^8x^5z^4 }}\times \sqrt{\frac{ x^9z^4 }{ 3 }}\] it is ^10 not O

OpenStudy (anonymous):

@sorour help please I have another one

OpenStudy (anonymous):

ok,i'll

OpenStudy (anonymous):